The f:[ 3,13] to R defined by f(x)= [ x-3]- 13 is continuous function. True or false with full explanation

1
Expert's answer
2022-03-18T01:28:35-0400

We have

"f(4-0)=0-13=-13\\
f(4)=1-13=-12\\
f(4-0)\ne f(4)"

Thus f is not continuous in x=4, hence it is not continuous on [3,13]